OpenStudy (anonymous):

solve 3 (sqrt) x + 4 - 2 = 13

OpenStudy (anonymous):

(sqrt) x +4 only

OpenStudy (anonymous):

\[\sqrt{x+4}=15\] \[x+4=15^{2}\] \[x=221\]